# presep= ]e1,e2,e3

### v1

@@ A.4 Proof of Theorem 4 (Functional CLT)
$\widetilde{\mathbb{T}}^{x}_{n}(s)$ $=\sqrt{k_{n}h_{n}}\left\{g(x)-g_{n}(x)\right\}\frac{\widetilde{T}^{x}_{n}(s)}{g(x)}+\frac{1}{g(x)}\widetilde{\mathbb{V}}^{x}_{n}(s)+\sqrt{k_{n}h_{n}}\frac{1}{g(x)}\left\{\mathbb{E}\left[\widetilde{V}^{x}_{n}(s)\right]-g(x)s^{-\alpha_{x}}\right\}.$
@@
[…]
Recall the definitions $u_{n}=F^{\leftarrow,x}\left(1-k_{n}/n\right)$ and $q_{n}=F^{\leftarrow,x}_{n}\left(1-k_{n}/n\right)$. Then $n\overline{F}^{x}(u_{n})=k_{n}$. Define the conditional tail empirical quantile function $\widetilde{Q}_{n}^{x}$ on $\left(0,k_{n}/n\right)$ by
$\widetilde{Q}_{n}^{x}(t)=\widetilde{T}_{n}^{\leftarrow,x}\left(\frac{\left\lfloor k_{n}t\right\rfloor}{k_{n}}\right)=\frac{q_{\left\lfloor k_{n}t\right\rfloor}}{u_{n}},$
and let $Q^{x}(t)=t^{-\frac{1}{\alpha_{x}}},t>0$.
@@
[…]
in $\mathbb{D}\left([s_{0},\infty)\right)\times\mathbb{D}((0,s_{0}^{-\frac{1}{\alpha_{x}}}])$ endowed with the product $J_{1}$ topology.
@@
[…]
@@
Note that $[s_{0},\infty)\ni s\mapsto T^{x}(s)$ is surjective on $(0,s_{0}^{-\frac{1}{\alpha_{x}}}]$, continuous and decreasing with negative derivative. As the processes $\widetilde{T}_{n}^{x}$ have non-increasing sample paths for each $n\geq 1$, Vervaat’s Lemma 20 yields that
[…]
in $\mathbb{D}\left([s_{0},\infty)\right)\times\mathbb{D}((0,s_{0}^{-\frac{1}{\alpha_{x}}}])$ endowed with the product $J_{1}$ topology. This concludes the proof. ∎
@@
$\text{var}\left(\frac{1}{\alpha_{x}}\mathbb{T}^{x}(1)\right)=\frac{1}{\alpha_{x}^{2}}\text{cov}(\mathbb{T}^{x}(1),\mathbb{T}^{x}(1))=\frac{1}{\alpha_{x}^{2}g(x)}\int K^{2}(u)\mathrm{d}u=\frac{\gamma(x)^{2}}{g(x)}\int K^{2}(u)\mathrm{d}u.$
@@
$\widehat{\mathbb{T}}^{x}_{n}=\widetilde{\mathbb{T}}^{x}_{n}\left(s\frac{q_{n}}{u_{n}}\right)+\sqrt{k_{n}h_{n}}\left\{T^{x}\left(s\frac{q_{n}}{u_{n}}\right)-T^{x}(s)\right\}.$
Recall that $\beta$-mixing is equivalent to being $m$-dependent. By Theorem 1, the tail consistency (4) holds, while Proposition 1 yields that $q_{n}/u_{n}\overset{\mathrm{Pr}}{\to}1$, and the continuous mapping theorem thus yields that the first term above converges to $\mathbb{T}^{x}$. As for the second term, Theorem 12 and the delta-method yields that
$\sqrt{k_{n}h_{n}}\left\{T^{x}\left(s\frac{q_{n}}{u_{n}}\right)-T^{x}(s)\right\}$ $=s^{-{\alpha_{x}}}\sqrt{k_{n}h_{n}}\left\{T^{x}\left(\frac{q_{n}}{u_{n}}\right)-T^{x}(1)\right\}$
[…]
jointly with $\widetilde{\mathbb{T}}^{x}_{n}\left(sq_{n}/u_{n}\right)$. The argument can be transferred unchanged to pathwise weak convergence. ∎
@@ A.5 Proof of Theorem 5 (Hill CLT)
$\sqrt{k_{n}h_{n}}\left(\widehat{\gamma}_{n}(x)-\gamma(x)\right)$ $=\sqrt{k_{n}h_{n}}\int_{q_{n}/u_{n}}^{1}\frac{\widetilde{T}_{n}^{x}\left(s\right)}{s}\mathrm{d}s+\int_{1}^{\infty}\frac{\widetilde{\mathbb{T}}_{n}^{x}\left(s\right)}{s}\mathrm{d}s.$
We examine the first term. Since $q_{n}/u_{n}\overset{\mathrm{Pr}}{\to}1$, we may assume without loss of generality that for given $\varepsilon\in(0,1)$, there exists $N\in\mathbb{N}$ such that for $n\geq N$ it holds that $q_{n}/u_{n}\in(1-\varepsilon,1+\varepsilon)$.
@@
$\sqrt{k_{n}h_{n}}\int_{q_{n}/u_{n}}^{1}\frac{\widetilde{T}_{n}^{x}\left(s\right)-T^{x}\left(s\right)}{s}\mathrm{d}s\leq\frac{\sqrt{k_{n}h_{n}}}{1-\varepsilon}\left|\frac{q_{n}}{u_{n}}-1\right|\sup_{t\in(1-\varepsilon,1+\varepsilon)}\left|\widetilde{T}_{n}^{x}\left(t\right)-T^{x}\left(t\right)\right|.$
@@
[…]
@@
$\sqrt{k_{n}h_{n}}\int_{q_{n}/u_{n}}^{1}\frac{\widetilde{T}_{n}^{x}\left(s\right)}{s}\mathrm{d}s=\sqrt{k_{n}h_{n}}\int_{q_{n}/u_{n}}^{1}\frac{T^{x}\left(s\right)}{s}\mathrm{d}s+o_{\mathrm{Pr}}(1).$
Applying the delta-method with $x\mapsto\int_{x}^{1}\frac{T^{x}\left(s\right)}{s}\mathrm{d}s$, on the sequence $(q_{n}/u_{n})$ yields that
[…]
@@
$\qquad+\sqrt{k_{n}h_{n}}\frac{1}{g(x)}\int_{1}^{\infty}\frac{\mathbb{E}\left[\widetilde{V}^{x}_{n}(s)\right]-g(x)s^{-\alpha_{x}}}{s}\mathrm{d}s.$
The last term is negligible due to condition 7. In the proof of Theorem 3 we showed $\int_{1}^{\infty}\widetilde{T}^{x}_{n}(s)/s\ \mathrm{d}s\overset{\mathrm{Pr}}{\to}\gamma(x)$. Combined with condition 4 we have that the first term also vanishes. Finally we consider $g(x)^{-1}\int_{1}^{\infty}\mathbb{V}_{n}^{x}\left(s\right)/s\ \mathrm{d}s$. It holds that
[…]
@@
$\text{var}\left(\mathbb{J}^{{\dagger}}_{n,A}\right)=\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n})}\text{var}\left(\sum_{j=1}^{r_{n}}K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\right)$
$\leq\frac{2m_{n}r_{n}}{nh_{n}\overline{F}^{x}(u_{n})}\sum_{j=0}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\right]$
@@
$\frac{m_{n}r_{n}}{nh_{n}\overline{F}^{x}(u_{n})}\sum_{j=0}^{m-1}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\right]$
$\leq\frac{m_{n}r_{n}}{nh_{n}\overline{F}^{x}(u_{n})}m\mathbb{E}\left[K^{2}\left(\frac{x-X_{0}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)^{2}\right].$
Then, recalling $m_{n}r_{n}=\mathcal{O}(n)$, and that the kernel $K$ has support $[-1,1]$, we have that $1_{\{(z-x)/h_{n}\leq 1\}}=1_{\{z\leq x+h_{n}\}}$ and hence we have the further bound
$\frac{m}{h_{n}\overline{F}^{x}(u_{n})}\int K^{2}\left(\frac{x-z}{h_{n}}\right)\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)^{2}\Big{|}X_{0}=z\right]g(z)\mathrm{d}z$
$\leq m\sup_{u\in[-h_{n}+x,x+h_{n}]}\left(\frac{1}{\overline{F}^{u}(u_{n})}\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)^{2}\Big{|}X_{0}=u\right]\right)\times\frac{1}{h_{n}}\int K^{2}\left(\frac{x-z}{h_{n}}\right)g(z)\frac{\overline{F}^{z}(u_{n})}{\overline{F}^{x}(u_{n})}\mathrm{d}z,$
@@
$\frac{1}{h_{n}}\int K^{2}\left(\frac{x-z}{h_{n}}\right)g(z)\frac{\overline{F}^{z}(u_{n})}{\overline{F}^{x}(u_{n})}\mathrm{d}z\leq\frac{1}{h_{n}}\int K^{2}\left(\frac{x-z}{h_{n}}\right)g(z)u_{n}^{-(c_{\alpha}+c_{L})\lvert x-z\rvert}\mathrm{d}z$
$\leq u_{n}^{(c_{\alpha}+c_{L})h_{n}}\int K^{2}\left(w\right)g(x-wh_{n})\mathrm{d}z=\mathcal{O}(1),$
@@
$\lim_{n\to\infty}\sup_{u\in[-h_{n}+x,x+h_{n}]}\left(\frac{1}{\overline{F}^{u}(u_{n})}\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)^{2}\Big{|}X_{0}=u\right]\right)$
$\leq\lim_{n\to\infty}\left(\frac{1}{\overline{F}^{x}(u_{n})}\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)^{2}\Big{|}X_{0}=x\right]\right)$
$\quad+\lim_{n\to\infty}\sup_{u\in[-h_{n}+x,x+h_{n}]}\Bigg{|}\frac{1}{\overline{F}^{u}(u_{n})}\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)^{2}\Big{|}X_{0}=u\right]-\frac{1}{\overline{F}^{x}(u_{n})}\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)^{2}\Big{|}X_{0}=x\right]\Bigg{|}$
$=\int_{A}^{\infty}\log^{2}(t)\alpha_{x}t^{-\alpha_{x}-1}\mathrm{d}t,$
where we have used Potter’s bound to invoke the dominated convergence theorem for the first limit, and the term in the difference can be bounded using Potter’s bound and applying twice the Lipschitz conditions on $L^{x}$ and $\alpha_{x}$ as was done in Theorem 3. The remaining upper sum is is asymptotically negligible by condition 6, i.e.
$\lim_{m\to\infty}\limsup_{n\to\infty}\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\right]}{h_{n}\overline{F}^{x}(u_{n})}=0.$
@@
[…]

### v2

@@ A.4 Proof of Theorem 4 (Functional CLT)
$\widetilde{\mathbb{T}}^{x}_{n}(s)$ $=\sqrt{k_{n}h_{n}}\left\{g(x)-g_{n}(x)\right\}\frac{\widetilde{T}^{x}_{n}(s)}{g(x)}+\frac{1}{g(x)}\widetilde{\mathbb{V}}^{x}_{n}(s)+\sqrt{k_{n}h_{n}}\frac{1}{g(x)}\left\{\mathbb{E}\left[\widetilde{V}^{x}_{n}(s)\right]-g(x)s^{-1/\gamma(x)}\right\}.$
@@
$\widetilde{\mathbb{T}}^{x}_{n}(s)=\sqrt{k_{n}h_{n}}\left\{g(x)-g_{n}(x)\right\}\frac{\widetilde{T}^{x}_{n}(s)}{g(x)}+\frac{1}{g(x)}\widetilde{\mathbb{V}}^{x}_{n}(s)+o(1)\overset{w}{\Longrightarrow}\frac{1}{g(x)}\mathbb{V}^{x}(s).$
Recall the definitions $u_{n}^{x}=F^{\leftarrow,x}\left(1-k_{n}/n\right)$ and $q_{n,k_{n}}^{x}=F^{\leftarrow,x}_{n}\left(1-k_{n}/n\right)$. Then $n\overline{F}^{x}(u_{n}^{x})=k_{n}$. Define the conditional tail empirical quantile function $\widetilde{Q}_{n}^{x}$ on $\left(0,k_{n}/n\right)$ by
$\widetilde{Q}_{n}^{x}(t)=\widetilde{T}_{n}^{\leftarrow,x}\left(\frac{\left\lfloor k_{n}t\right\rfloor}{k_{n}}\right)=\frac{q_{\left\lfloor k_{n}t\right\rfloor,n}}{u_{n}^{x}},$
and let $Q^{x}(t)=t^{-\gamma(x)},t>0$.
@@
$\sqrt{h_{n}k_{n}}\left\{\widetilde{T}_{n}^{x}-T^{x},\widetilde{Q}_{n}^{x}-Q^{x}\right\}\overset{w}{\Longrightarrow}\left(\mathbb{T}^{x},-\left(Q^{x}\right)^{\prime}\cdot\mathbb{T}^{x}\circ Q^{x}\right),$
in $\mathbb{D}\left([s_{0},\infty)\right)\times\mathbb{D}((0,s_{0}^{-\gamma(x)}])$ endowed with the product $J_{1}$ topology.
@@
$\sqrt{h_{n}k_{n}}\left\{\frac{q_{n,k_{n}}^{x}}{u_{n}^{x}}-1\right\}\overset{d}{\to}\gamma(x)\mathbb{T}^{x}(1),$
@@
Note that $s_{0},\infty)\ni s\mapsto T^{x}(s)$ is surjective on $(0,s_{0}^{-\gamma(x)}]$, continuous and decreasing with negative derivative. As the processes $\widetilde{T}_{n}^{x}$ have non-increasing sample paths for each $n\geq 1$, Vervaat’s Lemma 20 yields that
$\sqrt{h_{n}k_{n}}\left\{\widetilde{T}_{n}^{x}-T^{x},\widetilde{T}_{n}^{\leftarrow,x}-T^{\leftarrow,x}\right\}\overset{w}{\Longrightarrow}\left(\mathbb{T}^{x},-\left(T^{\leftarrow,x}\right)^{\prime}\cdot\mathbb{T}^{x}\circ T^{\leftarrow,x}\right),$
in $\mathbb{D}\left([s_{0},\infty)\right)\times\mathbb{D}((0,s_{0}^{-\gamma(x)}])$ endowed with the product $J_{1}$ topology. This concludes the proof. ∎
@@
$\text{var}\left(\gamma(x)\mathbb{T}^{x}(1)\right)=\gamma(x)^{2}\text{cov}(\mathbb{T}^{x}(1),\mathbb{T}^{x}(1))=\frac{\gamma(x)^{2}}{g(x)}\int K^{2}(u)\mathrm{d}u.$
@@
$\widehat{\mathbb{T}}^{x}_{n}=\widetilde{\mathbb{T}}^{x}_{n}\left(s\frac{q_{n,k_{n}}^{x}}{u_{n}^{x}}\right)+\sqrt{k_{n}h_{n}}\left\{T^{x}\left(s\frac{q_{n,k_{n}}^{x}}{u_{n}^{x}}\right)-T^{x}(s)\right\}.$
The continuous mapping theorem and Theorem 12 yields that the first term above converges to $\mathbb{T}^{x}$. As for the second term, Theorem 12 and the delta-method yields that
$\sqrt{k_{n}h_{n}}\left\{T^{x}\left(s\frac{q_{n,k_{n}}^{x}}{u_{n}^{x}}\right)-T^{x}(s)\right\}$ $=s^{-{1/\gamma(x)}}\sqrt{k_{n}h_{n}}\left\{T^{x}\left(\frac{q_{n,k_{n}}^{x}}{u_{n}^{x}}\right)-T^{x}(1)\right\}$
$\overset{d}{\to}s^{-{1/\gamma(x)}}\left(T^{x}\right)^{\prime}(1)\cdot\gamma(x)\mathbb{T}^{x}(1)=-T^{x}(s)\mathbb{T}^{x}(1),$
jointly with $\widetilde{\mathbb{T}}^{x}_{n}\left(sq_{n,k_{n}}^{x}/u_{n}^{x}\right)$. The argument can be transferred unchanged to pathwise weak convergence. ∎
@@ A.5 Proof of Theorem 5 (Hill CLT)
$\sqrt{k_{n}h_{n}}\left(\widehat{\gamma}_{n}(x)-\gamma(x)\right)$ $=\sqrt{k_{n}h_{n}}\int_{q_{n,k_{n}}^{x}/u_{n}^{x}}^{1}\frac{\widetilde{T}_{n}^{x}\left(s\right)}{s}\mathrm{d}s+\int_{1}^{\infty}\frac{\widetilde{\mathbb{T}}_{n}^{x}\left(s\right)}{s}\mathrm{d}s.$
We examine the first term. Since Theorem 12 yields $q_{n,k_{n}}^{x}/u_{n}^{x}\overset{\mathrm{Pr}}{\to}1$, we may assume without loss of generality that for given $\varepsilon\in(0,1)$, there exists $N\in\mathbb{N}$ such that for $n\geq N$ it holds that $q_{n,k_{n}}^{x}/u_{n}^{x}\in(1-\varepsilon,1+\varepsilon)$.
@@
$\sqrt{k_{n}h_{n}}\left|\int_{q_{n,k_{n}}^{x}/u_{n}^{x}}^{1}\frac{\widetilde{T}_{n}^{x}\left(s\right)-T^{x}\left(s\right)}{s}\mathrm{d}s\right|\leq\frac{\sqrt{k_{n}h_{n}}}{1-\varepsilon}\left|\frac{q_{n,k_{n}}^{x}}{u_{n}^{x}}-1\right|\sup_{t\in(1-\varepsilon,1+\varepsilon)}\left|\widetilde{T}_{n}^{x}\left(t\right)-T^{x}\left(t\right)\right|.$
@@
$\sqrt{k_{n}h_{n}}\left|\frac{q_{n,k_{n}}^{x}}{u_{n}^{x}}-1\right|\overset{d}{\to}\gamma(x)\left|\mathbb{T}^{x}(1)\right|,$
@@
$\sqrt{k_{n}h_{n}}\int_{q_{n,k_{n}}^{x}/u_{n}^{x}}^{1}\frac{\widetilde{T}_{n}^{x}\left(s\right)}{s}\mathrm{d}s=\sqrt{k_{n}h_{n}}\int_{q_{n,k_{n}}^{x}/u_{n}^{x}}^{1}\frac{T^{x}\left(s\right)}{s}\mathrm{d}s+o_{\mathrm{Pr}}(1).$
Applying the delta-method with $x\mapsto\int_{x}^{1}\frac{T^{x}\left(s\right)}{s}\mathrm{d}s$, on the sequence $(q_{n,k_{n}}^{x}/u_{n}^{x})$ yields that
$\sqrt{k_{n}h_{n}}\int_{q_{n,k_{n}}^{x}/u_{n}^{x}}^{1}\frac{T^{x}\left(s\right)}{s}\mathrm{d}s\overset{d}{\to}-\gamma(x)\mathbb{T}^{x}(1).$
@@
$\qquad+\sqrt{k_{n}h_{n}}\frac{1}{g(x)}\int_{1}^{\infty}\frac{\mathbb{E}\left[\widetilde{V}^{x}_{n}(s)\right]-g(x)s^{-1/\gamma(x)}}{s}\mathrm{d}s.$
The last term is negligible due to condition 7. Theorem 12 yields that equation (2) holds, and consequently $\int_{1}^{\infty}\widetilde{T}^{x}_{n}(s)/s\ \mathrm{d}s\overset{\mathrm{Pr}}{\to}\gamma(x)$ as in the proof of Theorem 3. Combined with condition 4 we have that the first term also vanishes. Finally we consider $g(x)^{-1}\int_{1}^{\infty}\widetilde{\mathbb{V}}_{n}^{x}\left(s\right)/s\ \mathrm{d}s$. It holds that
$\int_{1}^{A}\frac{\widetilde{\mathbb{V}}_{n}^{x}\left(s\right)}{s}\mathrm{d}s\underset{n\to\infty}{\overset{d}{\to}}\int_{1}^{A}\frac{\widetilde{\mathbb{V}}^{x}\left(s\right)}{s}\mathrm{d}s,\qquad\int_{A}^{\infty}\frac{\widetilde{\mathbb{V}}^{x}\left(s\right)}{s}\mathrm{d}s\underset{A\to\infty}{\overset{d}{\to}}0.$
@@
$\text{var}\left(\mathbb{J}^{{\dagger}}_{n,A}\right)=\frac{m_{n}}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\text{var}\left(\sum_{j=1}^{r_{n}}K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\right)$
$\leq\frac{2m_{n}r_{n}}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\sum_{j=0}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\right]$
@@
$\frac{m_{n}r_{n}}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\sum_{j=0}^{m-1}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\right]$
$\leq\frac{m_{n}r_{n}}{nh_{n}\overline{F}^{x}(u_{n}^{x})}m\mathbb{E}\left[K^{2}\left(\frac{x-X_{0}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)^{2}\right].$
Then, recalling $m_{n}r_{n}=\mathcal{O}(n)$, and that the kernel $K$ has support [1,1][-1,1], we have that $1_{\{(z-x)/h_{n}\leq 1\}}=1_{\{z\leq x+h_{n}\}}$ and hence we have the further bound
$\frac{m}{h_{n}\overline{F}^{x}(u_{n}^{x})}\int K^{2}\left(\frac{x-z}{h_{n}}\right)\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)^{2}\Big|X_{0}=z\right]g(z)\mathrm{d}z$
$\leq m\sup_{u\in[-h_{n}+x,x+h_{n}]}\left(\frac{1}{\overline{F}^{u}(u_{n}^{x})}\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)^{2}\Big|X_{0}=u\right]\right)\times\frac{1}{h_{n}}\int K^{2}\left(\frac{x-z}{h_{n}}\right)g(z)\frac{\overline{F}^{z}(u_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}\mathrm{d}z,$
@@
$\frac{1}{h_{n}}\int K^{2}\left(\frac{x-z}{h_{n}}\right)g(z)\frac{\overline{F}^{z}(u_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}\mathrm{d}z\leq\frac{1}{h_{n}}\int K^{2}\left(\frac{x-z}{h_{n}}\right)g(z)(u_{n}^{x})^{-(c_{\gamma}+c_{L})\lvert x-z\rvert}\mathrm{d}z$
$\leq(u_{n}^{x})^{(c_{\gamma}+c_{L})h_{n}}\int K^{2}\left(w\right)g(x-wh_{n})\mathrm{d}w=\mathcal{O}(1),$
@@
$\lim_{n\to\infty}\sup_{u\in[-h_{n}+x,x+h_{n}]}\left(\frac{1}{\overline{F}^{u}(u_{n}^{x})}\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)^{2}\Big|X_{0}=u\right]\right)$
$\leq\lim_{n\to\infty}\left(\frac{1}{\overline{F}^{x}(u_{n}^{x})}\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)^{2}\Big|X_{0}=x\right]\right)$
$\quad+\lim_{n\to\infty}\sup_{u\in[-h_{n}+x,x+h_{n}]}\Bigg|\frac{1}{\overline{F}^{u}(u_{n}^{x})}\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)^{2}\Big|X_{0}=u\right]-\frac{1}{\overline{F}^{x}(u_{n}^{x})}\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)^{2}\Big|X_{0}=x\right]\Bigg|$
$=\int_{A}^{\infty}\log^{2}(t)1/\gamma(x)t^{-1/\gamma(x)-1}\mathrm{d}t,$
where we have used Potter’s bound to invoke the dominated convergence theorem for the first limit, and the term in the difference can be bounded using Potter’s bound and applying twice the Lipschitz conditions on $L^{x}$ and $1/\gamma(x)$ as was done in Theorem 3. The remaining upper sum is is asymptotically negligible by condition 6, i.e.
$\lim_{m\to\infty}\limsup_{n\to\infty}\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}=0.$
@@
$\sqrt{k_{n}h_{n}}\left(\widehat{\gamma}_{n}\left(x\right)-\gamma(x)\right)\overset{d}{\to}\frac{1}{g(x)}\int_{1}^{\infty}\frac{\mathbb{V}^{x}\left(s\right)}{s}\mathrm{d}s-\frac{\gamma(x)}{g(x)}\mathbb{V}^{x}(1).$